Two students solved a problem involving a quadratic equation. The first student made an error only in the constant term of the equation and determined the roots were 2 and 8. The second student made an error only in the coefficient of the linear term and determined the roots were -1 and -9. What was the quadratic equation ?
Answer by ankor@dixie-net.com(22740)

Two students solved a problem involving a quadratic equation. The first student made an error only in the constant term of the equation and determined the roots were 2 and 8. The second student made an error only in the coefficient of the linear term and determined the roots were -1 and -9. What was the quadratic equation ?
:
Find the quadratic equation for each student's roots:
:
1st Student: (x-2)*(x-8) = x^2 - 10x + 16; Third term 16 is wrong
:
2nd Student: (x+1)*(x+9) = x^2 + 10x + 9; 2nd term +10 is wrong
:
I am going to say the correct equation is: x^2 - 10x + 9
:
Factors to (x-9)(x-1); Roots: x = 9, x = 1